Question 6 a) As a promotional offer, a new gym is offering a 30% discount to those who sign up in the first month. Patty and her 2 siblings take advantage of the offer and sign up for total amount of £210. What were the total savings made? Show your working.


1
Expert's answer
2021-10-21T14:26:29-0400

Discount=30%

Let amount per person=$x

Discounted amount per person

"=x-\frac{30}{100}x"

"=0.7x"

Total amount for three people

"=3(0.7x)=210\\x=100"

Total Savings

"=3x-210\\=3(100)-210=90"

Total savings is $90